OpenStudy (anonymous):

Verify: 1/cscX+1 + 1/csc-1 = 2tanX/cos Should I find the common denominator and then simplify? I keep trying to do this but end up at cscX-1/csc^2X-1 + cscX+1/csc^2X It looks all wrong. Please help. T_T

OpenStudy (anonymous):

Sorry I left some of the X's out. Jut noticed. :x

OpenStudy (anonymous):

deal with cscx+1

OpenStudy (anonymous):

\[\csc x+1=\frac{ 1+\sin x }{ \sin x }\] \[\csc x-1=\frac{ 1-\sin x }{ \sin x }\] inverted

OpenStudy (anonymous):

\[\frac{ \sin x }{\sin x +1}+\frac{ \sin x }{ \sin x-1 }\]

OpenStudy (anonymous):

Thank you, I'm going to try to work this out now from the beginning to make sure I am understanding this. I've never seen sec presented that way.

OpenStudy (anonymous):

\[\csc x=\frac{ 1 }{ \sin x }\] so add 1 both sides